OpenStudy (anonymous):

36x^6-63x^4+45x^3/-9x^2 Please simplify answer

OpenStudy (accessdenied):

\[ \begin{split} &\frac{36x^{6} - 63x^{4} + 45x^{3}}{-9x^{2}}\\ &\\ &=\frac{1}{-9x^{2}}(36x^{6} - 63x^{4} + 45x^{3})\\ &\\ &= \frac{36x^{6}}{-9x^{2}} + \frac{-63x^{4}}{-9x^{2}} + \frac{45x^{3}}{-9x^{2}}\\ \end{split} \] Are you able to simplify it at this point?

OpenStudy (accessdenied):

Oh, OP left. Well, I'll just leave the answer... \[ \begin{split} &\frac{36x^{6}}{-9x^{2}} + \frac{-63x^{4}}{-9x^{2}} + \frac{45x^{3}}{-9x^{2}}\\ &\\ &=\frac{36}{-9}\frac{x^{6}}{x^{2}} + \frac{-63}{-9}\frac{x^{4}}{x^{2}} + \frac{45}{-9}\frac{x^{3}}{x^{2}}\\ &\\ &=-4x^{6 -2} + 7x^{4 -2} - 5x^{3 -2}\\ &\\ &=-4x^{4} + 7x^{2} - 5x\\ \end{split} \]

OpenStudy (anonymous):

thank you so much for explaining it in detail

OpenStudy (accessdenied):

No problem! I hope that the process makes sense. :)